OpenStudy (madhatter):

plz help use full sentences If you know the number of yards for a measurement, you can change that measure to meters by multiplying the number of yards by 0.9144. a. Write a function rule that relates meters to yards. b. How many meters are equivalent to 7,200 yards? Round your answer to the nearest hundredth, if necessary. c. How many yards are equivalent to 2,500 meters? Round your answer to the nearest hundredth, if necessary.

OpenStudy (campbell_st):

(a) let M = metres and Y = Yards M = 0.9144Y (b) substitute Y = 7200 into the above function and round to 2 decimal places. \[M = 0.9144 \times 7200\] (c) you have an equation 2500 = 0.9144Y divide both sides of the equation by 0.9144 to find the number of yards. hope this helps

jimthompson5910 (jim_thompson5910):

Meters = 0.9144 times Yards M = 0.9144Y So that's how campbell_st got part a)

OpenStudy (madhatter):

oh ok

jimthompson5910 (jim_thompson5910):

if you want to write that in function notation, you would replace M with F(Y), f(Y) or something like that to get F(Y) = 0.9144Y

jimthompson5910 (jim_thompson5910):

and the reason why there aren't many steps here is because it literally says "you can change that measure to meters by multiplying the number of yards by 0.9144."

jimthompson5910 (jim_thompson5910):

so it's kinda given, which means not much work is to be done/shown

OpenStudy (madhatter):

why would it be that??

jimthompson5910 (jim_thompson5910):

part b) Plug in Y = 7200 to find M M = 0.9144Y M = 0.9144*7200 M = 6,583.68 So 6,583.68 meters are in 7200 yards

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (madhatter):

i mean y would it be F(Y)

jimthompson5910 (jim_thompson5910):

part c) plug in M = 2500 and solve for Y M = 0.9144Y 2500 = 0.9144Y 2500/0.9144 = Y 2,734.03324584427 = Y Y = 2,734.03324584427 Y = 2,734.03 So there are roughly 2,734.03 yards in 2500 meters

jimthompson5910 (jim_thompson5910):

because function notation is like f(x), p(x), etc the x is the input to the function in this case, the Y (yards) is the input which is why it's f(Y) or F(Y), etc

OpenStudy (madhatter):

oh ic lol

OpenStudy (madhatter):

thanks for the help

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (madhatter):

i just have 10 more problems to go

jimthompson5910 (jim_thompson5910):

fun fun...lol

OpenStudy (madhatter):

yep